# Robustness of quantum algorithms: Worst-case fidelity bounds and implications for design

### v1

@@ Abstract
Errors occurring on noisy hardware pose a key challenge to reliable quantum computing. Existing techniques such as error correction, mitigation, or suppression typically separate the error handling from the algorithm analysis and design. In this paper, we develop an alternative, algorithm-centered framework for understanding and improving the robustness against errors. For a given quantum algorithm and error model, we derive worst-case fidelity bounds which can be explicitly computed to certify the robustness. We consider general error models including coherent and (Markovian) incoherent errors and allowing for set-based error descriptions to address uncertainty or time-dependence in the errors. Our results give rise to guidelines for robust algorithm design and compilation by optimizing our theoretical robustness measure. Numerical results on algorithm analysis and robust optimization demonstrate the practicality of the framework.
preprint: APS/123-QED
I Introduction
@@
In [30], results in this direction were obtained in the form of worst-case fidelity bounds for the specific class of coherent control errors, which are multiplicative Hamiltonian errors that can be caused, e.g., by miscalibration. Analogous results for coherent control errors in quantum annealing were obtained in [31]. Further robustness analysis results include [32, 33], which study the average fidelity of individual gates in the presence of incoherent errors. Moreover, the recent paper [34] quantifies the average fidelity of an algorithm subject to noise using tools from quantum information geometry. They apply their theoretical results to derive resilience-runtime tradeoffs, showing, e.g., that longer algorithms can be more robust than shorter ones. However, [34] handles incoherent errors by averaging over coherent errors, which excludes relaxation errors such as amplitude damping. Our main contributions in the context of the above works are:
@@
    a flexible and general robustness analysis framework which can handle arbitrary coherent and (stationary or time-dependent Markovian) incoherent errors;
@@
    new insights into inherent robustness properties and fundamental limitations of quantum algorithms;
@@
The remainder of the paper is structured as follows. In Section II, we derive worst-case fidelity bounds for quantum algorithms subject to coherent errors. We develop explicit procedures for efficiently and accurately computing the bounds for a given algorithm and error model, and we discuss implications of our results for analysis and design. Next, in Section III, we provide analogous results for more general quantum algorithms (including invertible quantum operations) and error classes (including incoherent errors). In Section IV, we apply our theoretical findings in simulation to analyze the robustness of given quantum algorithms (composite pulses and the quantum Fourier transform) and to design or compile quantum algorithms with specified robustness properties. Finally, in Section V, we conclude the paper.
II Robustness of quantum algorithms against coherent errors
In this section, we analyze the robustness of quantum algorithms against coherent errors. Coherent errors are errors that can be described by unitary operations, and they pose serious challenges for reliably implementing quantum algorithms [42, 43, 44, 45]. After introducing the problem setup (Section II.1), we present the main result of this section: a worst-case fidelity bound (Section II.2). In Section II.3, we discuss the theoretical results, and we highlight their implications for quantum algorithm design in Section II.4.
II.1 Problem setup
@@
Uncertainty descriptions such as (4) are commonly referred to as set membership uncertainty, and they are well-established for modeling uncertainty, e.g., in estimation and control [7, 8, 9]. Note the generality of the above error model, which includes arbitrary unitary operators which can vary between qubits and gates. In particular, the set-based description includes independent errors which vary over time, e.g., between different executions of the algorithm, or between different hardware platforms, as long as the error bounds are fulfilled. In practice, estimates of the set $\mathcal{H}_{\mathrm{e},j}$ can be obtained, e.g., using tools for quantum process tomography [47] or error characterization [48, 49, 50].
@@
II.2 Worst-case fidelity bound
@@
$F_{\mathrm{wc}}\geq 1-\delta^{2}N^{2}\Big{(}\frac{N-1}{2}\delta+\gamma\Big{)}^{2}.$ (14)
The proof of Theorem II.1 can be found in Appendix A. Inequality (14) shows that the worst-case infidelity is bounded in terms of the error level $\delta$, the depth $N$, and the norm of the averaged interaction Hamiltonian $\lVert G\rVert$. Thus, the worst-case fidelity $F_{\mathrm{wc}}$ is high if the individual gate fidelity is high ($\delta$ is small) and the norm of the averaged interaction Hamiltonian $\lVert G\rVert$ is small. Note that, by definition, $G$ depends linearly on the error which is why the bound (13) is linear in $\delta$.
In Appendix B, we show how the bound $\gamma$ and, in particular, the worst-case fidelity bound (11) can be computed explicitly for a given quantum algorithm and error model. We present different complementary approaches with varying levels of scalability and conservatism, which can be scaled to large numbers of qubits and gates via a partitioning approach explained in Appendix C. Together with Theorem II.1, this yields a method for computing worst-case fidelity bounds for a given quantum algorithm $\bar{U}$ and an error description $\mathcal{U}_{\mathrm{e}}$. We note that a related result to Theorem II.1 was obtained in [52] in the context of robust quantum control. The main difference of our work is that we study discrete quantum algorithms whereas [52] considers (continuous-time) differential equations, which require different technical tools and admit partially more conservative results.
II.3 Discussion
Theorem II.1 implies that a given worst-case fidelity bound $\bar{F}_{\mathrm{wc}}$ is guaranteed if  [^3] The bound (15) uses N1<NN-1<N.
$\delta^{2}N^{2}\Big{(}\frac{\delta N}{2}+\gamma\Big{)}^{2}\leq 1-\bar{F}_{\mathrm{wc}}.$ (15)
Note that the noise level $\delta$ and the algorithm depth $N$ enter multiplicatively, i.e., when the algorithm depth $N$ increases then the worst-case fidelity necessarily deteriorates unless the noise level $\delta$ decreases proportionally, and vice versa. In Appendix D, we discuss why such a linear dependence $\delta\sim\frac{1}{N}$ can be expected for a worst-case fidelity bound. There, we also discuss the relation of the bound in Theorem II.1 to the existing literature, in particular to the robustness analysis for the special class of coherent control errors from [30] and to the average fidelity expression for stochastic noise models from [34].
The lower bound (15) has two main ingredients: The term $\gamma$ quantifies the contribution of the algorithm to the worst-case fidelity, which can be leveraged for robust algorithm design or compilation, compare Section II.4. On the other hand, even when $\gamma=0$, the bound
$\frac{1}{4}\delta^{4}N^{4}\leq 1-\bar{F}_{\mathrm{wc}},$ (16)
@@ II.3 Discussion
For larger values of $\gamma$, the required value of $\delta N$ decreases further. For example, for an algorithm with poor robustness, e.g., $\gamma=1$, achieving a worst-case fidelity bound of $\bar{F}_{\mathrm{wc}}\geq 0.999$ requires a smaller error level $\delta\leq 3.1\cdot 10^{-5}$ when keeping the depth at $N=10^{3}$. This shows that an algorithm which is designed in an inherently robust way, i.e., such that $\gamma\approx 0$, admits a significant increase of the tolerable error level $\delta$. In particular, for the fixed depth $N=10^{3}$ considered above, reducing $\gamma$ from $\gamma=1$ (low robustness) to $\gamma=0$ (high robustness) allows one to tolerate eight times larger error levels without reducing the guaranteed worst-case fidelity. Due to the linear dependence $\delta\sim\frac{1}{N}$, the required error levels $\delta$ for different depths $N$ can be obtained easily, e.g., increasing $N$ by a factor of $100$ decreases the allowable error level by a factor of $100$. In Section II.4, we discuss how to transfer these theoretical insights into concrete methods for (optimization-based) quantum algorithm design.
@@
It should be pointed out that Theorem II.1 only provides a lower bound on the worst-case fidelity and, therefore, the actual worst-case fidelity obtained by the algorithm can be larger. In the numerical simulations below, we show that the bound is a good qualitative indication of robustness, i.e., it directly correlates with empirical fidelity estimates. Further reducing conservatism to obtain more accurate robustness measures is an interesting direction for future research.
@@
Finally, let us discuss the influence of the number of qubits on the bound in Theorem II.1. To this end, note that multiple errors acting on $\ell$ of the qubits can be addressed via $H_{\mathrm{e},j}=\sum_{k=1}^{\ell}\theta_{j,k}B_{j,k}$ with $\lVert B_{j,k}\rVert\leq 1$ and unknown but bounded error angles $\theta_{j}=(\theta_{j,1},\dots,\theta_{j,k})$. In that case, (12) is guaranteed as long as $\lVert\theta_{j}\rVert\leq\delta$, $j=1,\dots,N$. Thus, for errors $\theta_{j}$ which are bounded in the $2$-norm, the worst-case fidelity bound (14) is independent of the number of qubits. Alternatively, an infinity-norm bound on $\theta_{j}$ such as $|\theta_{j,k}|\leq\tilde{\delta}$ for $j=1,\dots,N$, $k=1,\dots,\ell$ implies that (12) holds with $\delta=\tilde{\delta}\ell$. Hence, for errors $\theta_{j}$ bounded in the $\infty$-norm, the worst-case fidelity guaranteed via Theorem II.1 scales as $\mathcal{O}(\ell^{4})$ when $\ell$ of the qubits are affected by individual errors.
II.4 Implications for quantum algorithm design
@@
In this section, we discuss the implications of Theorem II.1 on quantum algorithm design and compilation. The result shows that the norm of $G$ can be used as a simple quantification of robustness of an algorithm against a given error model. In particular, according to (14), smaller values of the bound $\gamma$ in (13) lead to a larger guaranteed worst-case fidelity. Hence, circuits with smaller $\gamma$ admit a higher degree of guaranteed robustness against the specified error model $\mathcal{U}_{\mathrm{e}}$ and can, therefore, be implemented more reliably on noisy quantum hardware. This motivates taking the minimization of $\gamma$ as an additional objective into account during algorithm design or when mapping a given algorithm to the hardware via circuit compilation.
@@
Consider now the following multi-objective optimization problem: Minimizing the cost function $f(\eta)$ as well as maximizing the resilience of the algorithm against errors. For parameterized quantum circuits as in (17), the corresponding averaged interaction Hamiltonian $G(\eta)$ in (10) as well as its bound $\gamma(\eta)$ in (13) both depend on the parameters $\eta$. Thus, the above-mentioned multi-objective problem can be formulated as minimizing both the cost $f(\eta)$ as well as the robustness measure $\gamma(\eta)$. In practice, this optimization problem can be handled either by minimizing a weighted sum of the two terms
@@
minimizes the cost $f(\eta)$ while guaranteeing a specified robustness that is quantified via the parameter $c>0$. In Section IV.1, we apply this idea for designing composite pulse sequences with specified robustness properties. One can also define multiple robustness measures $\gamma_{l}(\eta)$, each of which corresponds to a different error class (e.g., Pauli X/Y/Z rotations, coherent control errors), and minimize a weighted sum
@@
III Robustness of quantum operations against coherent and incoherent errors
In this section, we generalize the framework from Section II by modeling circuit elements as well as errors as quantum operations. This allows us to consider not only more general (possibly non-unitary) circuit operations but also more general classes of errors including both coherent and (Markovian) incoherent errors. After introducing the problem setup in Section III.1, we present our main result in the form of a worst-case fidelity bound (Section III.2).
III.1 Problem setup
@@
$\mathcal{A}$ $=\bar{\mathcal{A}}_{N}\mathcal{A}_{\mathrm{e},N}\cdots\bar{\mathcal{A}}_{1}\mathcal{A}_{\mathrm{e},1}$ (27)
@@
As in Section II, the precise form of the quantum operation defined via (29) is unknown, i.e., the matrices $M_{\mathrm{e},j}$ are unknown. Instead, we assume that an uncertainty set describing the vectorized errors $\mathcal{A}_{\mathrm{e},j}$ is available, i.e., a set $\hat{\mathcal{A}}_{\mathrm{e}}$ of the form
@@
III.2 Worst-case fidelity bound

### v2

@@ Abstract
Errors occurring on noisy hardware pose a key challenge to reliable quantum computing. Existing techniques such as error correction, mitigation, or suppression typically separate the error handling from the algorithm analysis and design. In this paper, we develop an alternative, algorithm-centered framework for understanding and improving the robustness against errors. For a given quantum algorithm and error model, we derive worst-case fidelity bounds which can be efficiently computed to certify the robustness. We consider general error models including coherent and (Markovian) incoherent errors and allowing for set-based error descriptions to address uncertainty or time-dependence in the errors. Our results give rise to guidelines for robust algorithm design and compilation by optimizing our theoretical robustness measure. We demonstrate the practicality of the framework with numerical results on algorithm analysis and robust optimization, including the robustness analysis of a 50-qubit modular adder circuit.
1 Introduction
@@
In [30], results in this direction were obtained in the form of worst-case fidelity bounds for the specific class of coherent control errors, which are multiplicative Hamiltonian errors that can be caused, e.g., by miscalibration. Analogous results for coherent control errors in quantum annealing were obtained in [31]. Further robustness analysis results include [32, 33], which study the average fidelity of individual gates in the presence of incoherent errors. Moreover, the recent paper [34] quantifies the average fidelity of an algorithm subject to noise using tools from quantum information geometry. They apply their theoretical results to derive resilience-runtime tradeoffs, showing, e.g., that longer algorithms can be more robust than shorter ones. However, [34] handles incoherent errors by averaging over coherent errors, which excludes relaxation errors such as amplitude damping. Further, the recent paper [35] employs metastability for scalable robustness analysis of quantum algorithms, assuming Pauli noise and without deriving explicit fidelity bounds. Finally, related robustness questions were studied in the context of quantum control [36, 37].
In the context of these works, our main contributions are:
@@
    a general and scalable robustness analysis framework which can handle arbitrary coherent and (stationary or time-dependent Markovian) incoherent errors;
@@
    quantitative insights into inherent robustness properties and fundamental limitations of quantum algorithms;
@@
The remainder of the paper is structured as follows. In Section 2, we derive worst-case fidelity bounds for quantum algorithms subject to coherent errors. We develop explicit procedures for efficiently and accurately computing the bounds for a given algorithm and error model, and we discuss implications of our results for analysis and design. Next, in Section 3, we provide analogous results for more general quantum algorithms (including invertible quantum operations) and error classes (including incoherent errors). In Section 4, we apply our theoretical findings in simulation to analyze the robustness of given quantum algorithms (composite pulses, the quantum Fourier transform, and a $50$-qubit modular adder circuit) and to design or compile quantum algorithms with specified robustness properties. Finally, in Section 5, we conclude the paper.
2 Robustness of quantum algorithms against coherent errors
In this section, we analyze the robustness of quantum algorithms against coherent errors. Coherent errors are errors that can be described by unitary operations, and they pose serious challenges for reliably implementing quantum algorithms [45, 46, 47, 48]. After introducing the problem setup (Section 2.1), we present the main result of this section: a worst-case fidelity bound (Section 2.2). In Section 2.3, we discuss the theoretical results, and we highlight their implications for quantum algorithm design in Section 2.4.
2.1 Problem setup
@@
Uncertainty descriptions such as (4) are commonly referred to as set membership uncertainty, and they are well-established for modeling uncertainty, e.g., in estimation and control [7, 8, 9]. Note the generality of the above error model, which includes arbitrary unitary operators which can vary between qubits and gates. In particular, the set-based description includes independent errors which vary over time, e.g., between different executions of the algorithm, or between different hardware platforms, as long as the error bounds are fulfilled. In practice, estimates of the set $\mathcal{H}_{\mathrm{e},j}$ can be obtained, e.g., using tools for quantum process tomography [49] or error characterization [50, 51, 52, 53].
@@
2.2 Worst-case fidelity bound
@@
$F_{\mathrm{wc}}\geq 1-\delta^{2}N^{2}\Big(\frac{N-1}{2}\delta+\gamma\Big)^{2}$ (14)
The proof of Theorem 2.1 can be found in Appendix A. Inequality (14) shows that the worst-case infidelity is bounded in terms of the error level $\delta$, the depth $N$, and the norm of the averaged interaction Hamiltonian $\lVert G\rVert$. Thus, the worst-case fidelity $F_{\mathrm{wc}}$ is high if the individual gate fidelity is high ($\delta$ is small) and the norm of the averaged interaction Hamiltonian $\lVert G\rVert$ is small. Note that, by definition, $G$ depends linearly on the error which is why the bound (13) is linear in $\delta$.
In Appendix B, we show how the bound $\gamma$ and, in particular, the worst-case fidelity bound (11) can be computed explicitly for a given quantum algorithm and error model. We present different complementary approaches with varying levels of scalability and conservatism, which can be scaled to large numbers of qubits and gates via a partitioning approach explained in Appendix C. Together with Theorem 2.1, this yields a method for computing worst-case fidelity bounds for a given quantum algorithm $\bar{U}$ and an error description $\mathcal{U}_{\mathrm{e}}$. We note that a related result to Theorem 2.1 was obtained in [37] in the context of robust quantum control.
2.3 Discussion
Theorem 2.1 implies that a given worst-case fidelity bound $\bar{F}_{\mathrm{wc}}$ is guaranteed if  [^3] The bound (15) uses N1<NN-1<N.
$1-\bar{F}_{\mathrm{wc}}\geq\delta^{2}N^{2}\Big(\frac{\delta N}{2}+\gamma\Big)^{2}.$ (15)
Note that the noise level $\delta$ and the algorithm depth $N$ enter multiplicatively, i.e., when the algorithm depth $N$ increases then the worst-case fidelity necessarily deteriorates unless the noise level $\delta$ decreases proportionally, and vice versa. In Appendix D, we discuss why such a linear dependence $\delta\sim\frac{1}{N}$ can be expected for a worst-case fidelity bound. There, we also discuss the relation of the bound in Theorem 2.1 to the existing literature, in particular to the robustness analysis for the special class of coherent control errors from [30] and to the average fidelity expression for stochastic noise models from [34].
The lower bound (15) has two main ingredients: The term $\gamma$ quantifies the contribution of the algorithm to the worst-case fidelity, which can be leveraged for robust algorithm design or compilation, compare Section 2.4. On the other hand, even when $\gamma=0$, the bound
$1-\bar{F}_{\mathrm{wc}}\geq\frac{1}{4}(\delta N)^{4},$ (16)
@@ 2.3 Discussion
For larger values of $\gamma$, the required value of $\delta N$ decreases further. For example, for an algorithm with poor robustness, e.g., $\gamma=1$, achieving a worst-case fidelity bound of $\bar{F}_{\mathrm{wc}}\geq 0.999$ requires a smaller error level $\delta\leq 3.1\cdot 10^{-5}$ when keeping the depth at $N=10^{3}$. This shows that an algorithm which is designed in an inherently robust way, i.e., such that $\gamma\approx 0$, admits a significant increase of the tolerable error level $\delta$. In particular, for the fixed depth $N=10^{3}$ considered above, reducing $\gamma$ from $\gamma=1$ (low robustness) to $\gamma=0$ (high robustness) allows one to tolerate eight times larger error levels without reducing the guaranteed worst-case fidelity. Due to the linear dependence $\delta\sim\frac{1}{N}$, the required error levels $\delta$ for different depths $N$ can be obtained easily, e.g., increasing $N$ by a factor of $100$ decreases the allowable error level by a factor of $100$. In Section 2.4, we discuss how to transfer these theoretical insights into concrete methods for (optimization-based) quantum algorithm design.
@@
It should be pointed out that Theorem 2.1 only provides a lower bound on the worst-case fidelity and, therefore, the actual worst-case fidelity obtained by the algorithm can be larger. In the numerical simulations below, we show that the bound is a good qualitative indication of robustness, i.e., it directly correlates with empirical fidelity estimates. Further reducing conservatism to obtain more accurate robustness measures is an interesting direction for future research.
@@
Next, let us discuss the influence of the number of qubits on the bound in Theorem 2.1. To this end, note that multiple errors acting on $\ell$ of the qubits can be addressed via $H_{\mathrm{e},j}=\sum_{k=1}^{\ell}\theta_{j,k}B_{j,k}$ with $\lVert B_{j,k}\rVert\leq 1$ and unknown but bounded error angles $\theta_{j}=(\theta_{j,1},\dots,\theta_{j,k})$. In that case, (12) is guaranteed as long as $\lVert\theta_{j}\rVert\leq\delta$, $j=1,\dots,N$. Thus, for errors $\theta_{j}$ which are bounded in the $2$-norm, the worst-case fidelity bound (14) is independent of the number of qubits. Alternatively, an infinity-norm bound on $\theta_{j}$ such as $|\theta_{j,k}|\leq\tilde{\delta}$ for $j=1,\dots,N$, $k=1,\dots,\ell$ implies that (12) holds with $\delta=\tilde{\delta}\ell$. Hence, for errors $\theta_{j}$ bounded in the $\infty$-norm, the worst-case fidelity guaranteed via Theorem 2.1 scales as $\mathcal{O}(\ell^{4})$ when $\ell$ of the qubits are affected by individual errors.
Finally, while this section focuses on unitary quantum algorithms, it can also include non-unitary effects via bipartite systems with system-bath interaction. To be precise, suppose that the ideal unitary gate $\bar{U}_{j}$ is given by
$\bar{U}_{j}=\bar{U}_{\mathrm{S},j}\otimes I_{\mathrm{B}},$ (17)
where $\bar{U}_{\mathrm{S},j}$ acts only on the main system and the identity $I_{\mathrm{B}}$ acts on the environment. Further, suppose that the error Hamiltonian defining the unitary error $U_{\mathrm{e},j}=e^{-iH_{\mathrm{e},j}}$ is equal to
$H_{\mathrm{e},j}=H_{\mathrm{e}\mathrm{S},j}\otimes I_{\mathrm{B}}+I_{\mathrm{S}}\otimes H_{\mathrm{e}\mathrm{B},j}+H_{\mathrm{e}\mathrm{S}\mathrm{B},j}$ (18)
with the identity $I_{\mathrm{S}}$ acting on the main system and error Hamiltonians for the system $H_{\mathrm{e}\mathrm{S},j}$, bath $H_{\mathrm{e}\mathrm{B},j}$, and system-bath interaction $H_{\mathrm{e}\mathrm{S}\mathrm{B},j}$. All results in this section equally apply to such bipartite systems which allows us to model incoherent effects as well. However, computing the fidelity bound in Theorem 2.1 requires us to keep track of the bath state without allowing for reset, which limits the applicability for practical noise models. Hence, in Section 3, we present an alternative and more flexible approach for handling decoherence by modeling the ideal algorithm and the errors as vectorized quantum operations.
2.4 Implications for quantum algorithm design
@@
In this section, we discuss the implications of Theorem 2.1 on quantum algorithm design and compilation. The result shows that the norm of $G$ can be used as a simple quantification of robustness of an algorithm against a given error model. In particular, according to (14), smaller values of the bound $\gamma$ in (13) lead to a larger guaranteed worst-case fidelity. Hence, circuits with smaller $\gamma$ admit a higher degree of guaranteed robustness against the specified error model $\mathcal{U}_{\mathrm{e}}$ and can, therefore, be implemented more reliably on noisy quantum hardware. This motivates taking the minimization of $\gamma$ as an additional objective into account during algorithm design or when mapping a given algorithm to the hardware via circuit compilation.
@@
Consider now the following multi-objective optimization problem: Minimizing the cost function $f(\eta)$ as well as maximizing the resilience of the algorithm against errors. For parameterized quantum circuits as in (19), the corresponding averaged interaction Hamiltonian $G(\eta)$ in (10) as well as its bound $\gamma(\eta)$ in (13) both depend on the parameters $\eta$. Thus, the above-mentioned multi-objective problem can be formulated as minimizing both the cost $f(\eta)$ as well as the robustness measure $\gamma(\eta)$. In practice, this multi-objective optimization problem can be handled either by minimizing a weighted sum of the two terms
@@
minimizes the cost $f(\eta)$ while guaranteeing a specified robustness that is quantified via the parameter $c>0$. In Section 4.1, we apply this idea for designing composite pulse sequences with specified robustness properties. One can also define multiple robustness measures $\gamma_{l}(\eta)$, each of which corresponds to a different error class (e.g., Pauli X/Y/Z rotations, coherent control errors), and minimize a weighted sum
@@
3 Robustness of quantum operations against coherent and incoherent errors
In this section, we generalize the framework from Section 2 by modeling circuit elements as well as errors as quantum operations. This allows us to consider not only more general (possibly non-unitary) circuit operations but also more general classes of errors including both coherent and (Markovian) incoherent errors. After introducing the problem setup in Section 3.1, we present our main result in the form of a worst-case fidelity bound (Section 3.2).
3.1 Problem setup
@@
$\mathcal{A}$ $=\bar{\mathcal{A}}_{N}\mathcal{A}_{\mathrm{e},N}\cdots\bar{\mathcal{A}}_{1}\mathcal{A}_{\mathrm{e},1},$ (29)
@@
As in Section 2, the precise form of the quantum operation defined via (31) is unknown, i.e., the matrices $M_{\mathrm{e},j}$ are unknown. Instead, we assume that an uncertainty set describing the vectorized errors $\mathcal{A}_{\mathrm{e},j}$ is available, i.e., a set $\hat{\mathcal{A}}_{\mathrm{e}}$ of the form
@@
3.2 Worst-case fidelity bound